# kNNSampler: Stochastic Imputations for Recovering Missing Value Distributions

### v1

@@ 4.1.1 Data Settings
This model, considered by  Lalande and Doya, (2023), randomly generates covariate $x_{i}$ and response $y_{i}$ for each unit $i=1,\dots,N$ from a noisy two-dimensional ring of unit radius perturbed with an additive Gaussian noise of variance $0.1$:
@@ 4.1.2 Performance Metric: Energy Distance
To quantify the performance of an imputation method in recovering the missing value distribution, we compute the energy distance (Székely and Rizzo,, 2013) between the empirical distributions of the complete and imputed datasets. The energy distance is a well-established distance metric between probability distributions and is a parameter-free special case of the MMD (Sejdinovic et al.,, 2013).
@@
This is an unbiased estimate of the squared energy distance between the two joint distributions $Q(x,y)=P(y|x)Q(x)$ and $Q^{*}(x,y)=P^{*}(y|x)Q(x)$, where $P(y|x)$ is the true conditional distribution of true response $y$ given covariate $x$, $P^{*}(y|x)$ is the conditional distribution of imputed response $y$ given covariate $x$, and $Q(x)$ is the covariate distribution of missing units.
@@
A lower energy distance means that the two joint distributions are more similar, implying a better recovery of the missing value distribution. A higher energy distance implies that the imputed distribution is more different from the true data distribution.
P-values
Along with the energy distance itself, we report the p-value from a permutation two-sample test using the energy distance as a test statistic, where the null hypothesis is the identity of the two joint distributions $Q(x,y)$ and $Q^{*}(x,y)$. A smaller p-value suggests more significantly that the null hypothesis is false, implying that the imputed distribution $Q^{*}(x,y)$ is dissimilar to the true missing value distribution $Q(x,y)$. A larger p-value suggests that there is no evidence of the null hypothesis being false, suggesting that the imputed dataset is not statistically distinguishable from the dataset with true missing responses. Therefore, a larger p-value implies a better distribution recovery performance of the imputation method.
@@ 4.1.3 Benchmark Imputation Methods
Random Forest (Stekhoven and Bühlmann,, 2012): This method, widely used in practice, imputes a missing response by averaging its multiple predictions made by bootstrap-sampled tree regressors. It can learn a nonlinear relation between the response and covariate and handle the interactions among covariate features (e.g., Shah et al.,, 2014; Tang and Ishwaran,, 2017). We use the default configuration in scikit-learn.
kNNImputer (Troyanskaya et al.,, 2001): See Section 2.2 for the description of the method. We set the number $k$ of nearest neighbours as $k=5$, which is the default setting in scikit-learn and widely used in practice.
@@ 4.2.2 Quantitative Comparisons
Each experiment, consisting of data generation, imputations by each method, and the calculation of an evaluation metric, was independently repeated 10 times, and the mean and standard deviation of the evaluation metric are reported.
Figures 5 and 6 report the results on the energy distance between the empirical distributions of the imputed and true missing values; Figures 10 and 11 in Appendix B show the results for the best two methods for easier comparison. Figures 7 and 8 report the corresponding p-values. See Section 4.1.2 for details.
Figure 5: The energy distance between the empirical distributions of imputations and true missing values for the linear chi-square data (15). For each method and sample size, the average and standard deviation over 10 independent runs are shown.
Figure 6: The energy distance between the empirical distributions of imputations and true missing values for the noisy ring data (16). For each method and sample size, the average and standard deviation over 10 independent runs are shown.
The energy distance for the linear imputer is the highest among the different methods, quantifying the large discrepancy between the distributions of the imputations and true missing values, as visually observed in Figures 3 and 4. In contrast, the root mean squared error (RMSE) for the linear imputer is the lowest (see Appendix 11), as the true regression function in either case is a linear function. This result demonstrates that the RMSE is not a good metric for evaluating the distributional similarity between imputations and missing values. See Näf et al., (2023) for a further discussion.
The energy distances for kNNImputer and Random Forest are lower than those of the linear imputer, but they are still significantly higher than those of the two other methods. This is reasonable because they are estimating the conditional mean of the missing response given a covariate.
@@
Figure 7: The p-value from a permutation two-sample test, the test statistic being the energy distance between the empirical distributions of imputed and true missing values, on the linear chi-square data (15). For each method and sample size, the average and standard deviation of the p-value over 10 independent runs are shown.
Figure 8: The p-value from a permutation two-sample test, the test statistic being the energy distance between the empirical distributions of imputed and true missing values, on the noisy ring data (16). For each method and sample size, the average and standard deviation of the p-value over 10 independent runs are shown.
@@
5 Conclusion
@@ Appendix A Proof of Theorem 1
We proceed as the proof of Kpotufe, (2011, Theorem 1) on real-valued kNN regression, with adaptations to our RKHS-valued kNN regression setting.
@@ A.1 Bias Bound
Lemma 1 below is from Kpotufe, (2011, Lemma 1).
@@
where $r_{n,k}(x)$ is the distance between $x$ and its $k$-th nearest neighbour in $X_{n}$. This distance is bounded as in the proof of Kpotufe, (2011, Lemma 2), which leads to the claimed bound. For completeness, we prove it here.
@@ A.2 Variance Bound
$=\left\|\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\left\{\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\}\right\|_{\mathcal{H}}\leq\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\left\|\left\{\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\}\right\|_{\mathcal{H}},$
where the third expression follows from the definition of $\Phi(\hat{P}(\cdot\mid x))$ in (8). The notation $\psi({\rm NN}(x,k,X_{n}))$ emphasizes that it depends only on the subset of training data $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ associated with the indices ${\rm NN}(x,k,X_{n})$ of the $k$-nearest neighbours of $x$ in $X_{n}=\{x_{1},\dots,x_{n}\}$.
Because of the bound (9), changing $y_{i}$ for any $i\in{\rm NN}(x,k,X_{n})$ to any different value $y^{\prime}_{i}\in\mathcal{Y}$ changes the value of $\psi({\rm NN}(x,k,X_{n}))$ at most $2\sqrt{2}C_{\rm ker}/k$, since
$\left|\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y_{i}}-\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}\right|$
$\leq\left|\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y_{i}}\right|+\left|\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}\right|$
$=\left.\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\left\|\left\{\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\}\right\|_{\mathcal{H}}\right|_{y_{i}}+\left.\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\left\|\left\{\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\}\right\|_{\mathcal{H}}\right|_{y_{i}^{\prime}}$
$=\frac{1}{k}\left(\left\|\left\{\Phi(P(\cdot\mid x_{i}))-\Phi(y_{i})\right\}\right\|_{\mathcal{H}}+\left\|\left\{\Phi(P(\cdot\mid x_{i}))-\Phi(y_{i}^{\prime})\right\}\right\|_{\mathcal{H}}\right)\leq\frac{2\sqrt{2}C_{\rm ker}}{k},$
where $\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}$ denotes that the value of $y_{i}$ in $\psi({\rm NN}(x,k,X_{n}))$ is replaced by $y^{\prime}_{i}$. On the other hand, the output $y_{i}$ associated with any non-$k$-nearest neighbours $i\not\in{\rm NN}(x,k,X_{n})$ does not appear in $\psi({\rm NN}(x,k,X_{n}))$, so changing the value of $y_{i}$ in this case does not change $\psi({\rm NN}(x,k,X_{n}))$.
@@
The number of distinct index sets of nearest neighbours is identical to the number of distinct ways the set $X_{n}$ of $n$ points is intersected by balls $B(x,r_{k,n}(x))$ of center $x$ and radius $r_{k,n}(x)$ being the distance of the $k$-th nearest neighbour from $x$. This number is upper-bounded by the number of distinct ways $X_{n}$ is intersected by the class $\mathcal{B}=\{B(x,r)\mid x\in\mathcal{X},\ r>0\}$ of all balls, which is further upper-bounded by $n^{\mathcal{V}_{\mathcal{B}}}$ with the VC dimension $\mathcal{V}_{\mathcal{B}}$ of $\mathcal{B}$ (Kpotufe,, 2011, p.6). Therefore, by using the union bound, the probability that the statement (23) holds for some $x$ and $k$ is upper bounded by the bound (22) times $n^{\mathcal{V}_{\mathcal{B}}}$:
@@
$\mathbb{E}\left[\left\|\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\left\{\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\}\right\|_{\mathcal{H}}^{2}\mid X_{n}\right]=\mathbb{E}\left[\left\|\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}z_{j}\right\|_{\mathcal{H}}^{2}\mid X_{n}\right]$
@@
Appendix B Additional Experimental Results
Figure 10: The energy distance for the best two methods from Figure 5; see its caption for details.
Figure 11: The energy distance for the best two methods from Figure 6; see its caption for details.
Figures 10 and 11 show the results on the energy distance for the three best methods in Figures 5 and 6, respectively. Figure 12 shows the corresponding results on the root mean squared errors (RMSEs) between the imputations and true missing values. As mentioned in the main body, having low RMSEs does not imply good imputations in terms of recovering the distribution of missing values. Indeed, imputations from the linear imputer have the lowest RMSEs, but their distribution significantly differs from the distribution of true missing values, as quantified in Figures 5 and 6 and visually observed in Figures 3 and 4.
Figure 12: Root mean squared error (RMSE) for each imputation method for different sample sizes. The mean and standard deviation over 10 independent runs are shown for each setting. The top and bottom figures are the results on the linear chi-square data and the noisy ring data, respectively.

### v2

@@ 4.1.1 Data Settings
This model, considered by  lalande2023numerical, randomly generates covariate $x_{i}$ and response $y_{i}$ for each unit $i=1,\dots,N$ from a noisy two-dimensional ring of unit radius perturbed with an additive Gaussian noise of variance $0.1$:
@@ 4.1.2 Performance Metric: Energy Distance
To quantify the performance of an imputation method in recovering the missing value distribution, we compute the energy distance (szekely2013energy) between the empirical distributions of the complete and imputed datasets. We use the energy distance as it is a proper distance between distributions, can be easily computed from samples based on their Euclidean distances without the need for optimization (as compared with, e.g., a Wasserstein distance whose computation requires optimization to solve optimal transport (peyre2019computational)), and is parameter-free (in contrast to the MMD defined with, e.g., a Gaussian kernel, which depends on the bandwidth parameter). The energy distance is a canonical instance of MMD defined with a distance-based kernel (sejdinovic2013equivalence): it is “canonical” in the sense that it is both scale-invariant (the distance scales linearly with the scale of data) and rotation-invariant (szekely2013energy, Section 3).
@@
This is an unbiased estimate of the squared energy distance between the two joint distributions $Q(x,y)=P(y|x)Q(x)$ and $Q^{*}(x,y)=P^{*}(y|x)Q(x)$, where $P(y|x)$ is the true conditional distribution of true response $y$ given covariate $x$, $P^{*}(y|x)$ is the conditional distribution of imputed response $y$ given covariate $x$, and $Q(x)$ is the covariate distribution of missing units:
@@
A lower energy distance indicates that the two joint distributions are more similar, implying better recovery of the missing-value distribution. A higher energy distance indicates that the imputed distribution is more dissimilar to the true data distribution.
@@ 4.1.3 Benchmark Imputation Methods
Random Forest (stekhoven2012missforest): This method, widely used in practice, imputes a missing response by averaging its multiple predictions made by bootstrap-sampled tree regressors. It can learn a nonlinear relation between the response and covariate and handle the interactions among covariate features (e.g., shah2014comparison; tang2017random). We use the default configuration in scikit-learn.
kNNImputer (troyanskaya2001missing): See Section 2.2 for the description of the method. We set the number $k$ of nearest neighbours as $k=5$, which is the default setting in scikit-learn and widely used in practice.
@@
As suggested earlier, the number $k$ of nearest neighbours for kNNSampler is determined by the fast leave-one-out cross-validation method of kanagawa2024fast using the observed covariate-response pairs.
@@ 4.2.2 Quantitative Comparisons
Each experiment, consisting of data generation, imputations by each method, and the calculation of the evaluation metric, was independently repeated 10 times, and the mean and standard deviation of the evaluation metric are reported. Tables 1 and 2 report the results on the energy distance between the empirical distributions of the imputed and true missing values. See Section 4.1.2 for details.
Table 1: The energy distance between the empirical distributions of imputations and true missing values on the linear chi-square dataset (15). For each method and sample size, the average and standard deviation over 10 independent runs are shown.
@@
3000 0.027 $\pm$ 0.031 0.076 $\pm$ 0.023 0.200 $\pm$ 0.038 0.036 $\pm$ 0.033 0.585 $\pm$ 0.053
5000 0.027 $\pm$ 0.009 0.077 $\pm$ 0.030 0.199 $\pm$ 0.041 0.033 $\pm$ 0.019 0.598 $\pm$ 0.025
7000 0.027 $\pm$ 0.021 0.080 $\pm$ 0.018 0.219 $\pm$ 0.024 0.028 $\pm$ 0.018 0.589 $\pm$ 0.034
9000 0.017 $\pm$ 0.009 0.076 $\pm$ 0.023 0.183 $\pm$ 0.031 0.016 $\pm$ 0.007 0.605 $\pm$ 0.054
11000 0.018 $\pm$ 0.011 0.080 $\pm$ 0.033 0.198 $\pm$ 0.040 0.026 $\pm$ 0.021 0.584 $\pm$ 0.034
Table 2: The energy distance between the empirical distributions of imputations and true missing values on the noisy ring dataset (16). For each method and sample size, the average and standard deviation over 10 independent runs are shown.
[…]
3000 0.021 $\pm$ 0.015 0.076 $\pm$ 0.025 0.181 $\pm$ 0.032 0.033 $\pm$ 0.017 0.584 $\pm$ 0.038
5000 0.019 $\pm$ 0.015 0.069 $\pm$ 0.023 0.216 $\pm$ 0.055 0.024 $\pm$ 0.013 0.576 $\pm$ 0.042
7000 0.028 $\pm$ 0.009 0.087 $\pm$ 0.031 0.189 $\pm$ 0.032 0.028 $\pm$ 0.015 0.612 $\pm$ 0.044
9000 0.028 $\pm$ 0.022 0.074 $\pm$ 0.027 0.197 $\pm$ 0.043 0.020 $\pm$ 0.013 0.593 $\pm$ 0.033
11000 0.019 $\pm$ 0.012 0.075 $\pm$ 0.027 0.194 $\pm$ 0.064 0.035 $\pm$ 0.040 0.606 $\pm$ 0.062
Table 3: The root mean squared error of each imputation method for different sample sizes on the linear chi-square dataset (15). The mean and standard deviation over 10 independent runs are shown for each setting.
[…]
3000 2.691 $\pm$ 0.151 2.338 $\pm$ 0.154 2.117 $\pm$ 0.158 2.876 $\pm$ 0.126 1.885 $\pm$ 0.195
5000 2.710 $\pm$ 0.134 2.273 $\pm$ 0.113 2.092 $\pm$ 0.123 2.726 $\pm$ 0.190 1.914 $\pm$ 0.088
7000 2.729 $\pm$ 0.102 2.307 $\pm$ 0.118 2.100 $\pm$ 0.185 2.789 $\pm$ 0.135 1.895 $\pm$ 0.121
9000 2.786 $\pm$ 0.228 2.308 $\pm$ 0.076 2.065 $\pm$ 0.097 2.812 $\pm$ 0.095 1.945 $\pm$ 0.076
11000 2.793 $\pm$ 0.188 2.388 $\pm$ 0.127 2.055 $\pm$ 0.116 2.708 $\pm$ 0.184 1.913 $\pm$ 0.154
Table 4: The root mean squared error of each imputation method for different sample sizes on the noisy ring dataset (16). The mean and standard deviation over 10 independent runs are shown for each setting.
[…]
3000 2.680 $\pm$ 0.238 2.309 $\pm$ 0.133 2.073 $\pm$ 0.169 2.811 $\pm$ 0.112 1.951 $\pm$ 0.108
5000 2.818 $\pm$ 0.195 2.322 $\pm$ 0.141 2.079 $\pm$ 0.157 2.698 $\pm$ 0.130 1.870 $\pm$ 0.123
7000 2.733 $\pm$ 0.216 2.307 $\pm$ 0.141 2.133 $\pm$ 0.163 2.799 $\pm$ 0.186 1.959 $\pm$ 0.179
9000 2.638 $\pm$ 0.146 2.281 $\pm$ 0.103 2.138 $\pm$ 0.141 2.637 $\pm$ 0.177 1.923 $\pm$ 0.157
11000 2.672 $\pm$ 0.137 2.281 $\pm$ 0.152 2.024 $\pm$ 0.075 2.763 $\pm$ 0.164 1.885 $\pm$ 0.152
[…]
For comparison, we also report the root mean squared error (RMSE) for each method’s imputations in Tables 3 and 4. RMSE is expected to be smaller for regression-based methods, which estimate the conditional means of missing values and thereby minimize RMSE. A smaller RMSE does not imply better recovery of the missing-value distribution. Indeed, imputations from the linear imputer have the lowest RMSEs, but their distribution significantly differs from the distribution of true missing values, as quantified in Figures 1 and 2 and visually observed in Figures 3 and 4. This result demonstrates that the RMSE is not a good metric for evaluating the distributional similarity between imputations and missing values. See naf2023imputation for a related discussion.
@@
5 Real Data Experiments
Lastly, we present real-data experiments on solar power generated by photovoltaic panels, where missing values are common due to sensor failures and other factors (e.g., phan2023enhancing; costa2024employing). We use a Kaggle dataset [^8] https://www.kaggle.com/datasets/samuelkamau/solar-data/ that contains solar panel DC powers (responses) and the corresponding irradiations (covariates), totaling 67,698 covariate-response pairs. We randomly select a subset of $N$ covariate-response pairs from the full dataset. In this subset, we select randomly $30\%$ of the units whose covariates are between 0.4 and 0.6 and set their responses to missing. These missing responses are imputed based on the remaining observed covariate-response pairs in the subset. We consider each of $N\in\{10,000,~20,000,~30,000,~40,000,~50,000,~60,000\}$. The configuration of each method follows Section 4.1.3.
[…]
Table 5: Comparison of the energy distance between the empirical distributions of imputations and true missing values across different sample sizes of the real solar panel dataset in Section 5. For each method and sample size, the average and standard deviation of the energy distance over 10 independent runs are shown.
[…]
10000 $\mathbf{1.855}\bm{\pm 1.474}$ $5.544\pm 1.916$ $10.619\pm 2.958$ $3.333\pm 2.609$ $190.546\pm 16.701$
20000 $\mathbf{0.687}\bm{\pm 0.671}$ $4.980\pm 0.977$ $7.389\pm 1.763$ $2.634\pm 1.238$ $195.623\pm 9.364$
30000 $\mathbf{0.500}\bm{\pm 0.309}$ $5.112\pm 1.207$ $4.874\pm 1.144$ $2.273\pm 0.990$ $195.412\pm 3.468$
40000 $\mathbf{0.373}\bm{\pm 0.317}$ $5.543\pm 0.684$ $4.584\pm 0.890$ $2.293\pm 1.165$ $194.081\pm 5.721$
50000 $\mathbf{0.190}\bm{\pm 0.138}$ $5.667\pm 0.968$ $4.161\pm 0.881$ $2.559\pm 1.133$ $198.261\pm 4.039$
60000 $\mathbf{0.148}\bm{\pm 0.077}$ $6.230\pm 0.812$ $4.634\pm 1.053$ $1.927\pm 0.942$ $194.808\pm 3.526$
6 Conclusion and Discussion
@@
We discuss limitations of the current work, some of which stem from hot-deck methods in general (see andridge2010review), and potential future directions. (i) While the experiments show the promising performance of kNNSampler, they are limited to low-dimensional covariates. Experiments with higher-dimensional covariates are needed to fully characterize the KNNSampler’s practical performance. (ii) For higher-dimensional covariates, the choice of the distance function itself influences kNNSampler’s performance, which should be investigated. (iii) Leave-one-out cross-validation for selecting the number of nearest neighbours uses the mean square error, which should be modified to a distributional error metric. (iv) A further theoretical analysis is needed to understand how the distributional imputation quality affects subsequent analysis of a quantity of interest, such as the well-calibratedness of uncertainty estimates.
Acknowledgments
We thank Frédéric Coutellier, Michele Bezzi and colleagues at SAP Labs France for their support and discussion. We also thank the reviewers and the Action Editor for their time and constructive feedback.
@@ Appendix A Proof of Theorem 1
We proceed as the proof of kpotufe2011k on real-valued kNN regression, with adaptations to our RKHS-valued kNN regression setting.
@@ A.1 Bias Bound
Lemma 1 below is from kpotufe2011k.
@@
where $r_{n,k}(x)$ is the distance between $x$ and its $k$-th nearest neighbour in $X_{n}$. This distance is bounded as in the proof of kpotufe2011k, which leads to the claimed bound. For completeness, we prove it here.
@@ A.2 Variance Bound
$=\left\|\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\|_{\mathcal{H}},$
where the last expression follows from the definition of $\Phi(\hat{P}(\cdot\mid x))$ in (8). The notation $\psi({\rm NN}(x,k,X_{n}))$ emphasizes that it depends only on the subset of training data $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ associated with the indices ${\rm NN}(x,k,X_{n})$ of the $k$-nearest neighbours of $x$ in $X_{n}=\{x_{1},\dots,x_{n}\}$.
Because of the bound (9), changing $y_{i}$ for any $i\in{\rm NN}(x,k,X_{n})$ to any different value $y^{\prime}_{i}\in\mathcal{Y}$ changes the value of $\psi({\rm NN}(x,k,X_{n}))$ at most $2\sqrt{2}C_{\rm ker}/k$. This can be shown as follows. Let us write the last expression of (21) with the original $y_{i}$ and the one with $y_{i}$ replaced by $y^{\prime}_{i}$ as
$\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y_{i}}=\left\|A+B\right\|_{\mathcal{H}},\quad\quad\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}=\left\|A^{\prime}+B\right\|_{\mathcal{H}},$
where
$A:=\frac{1}{k}\left(\Phi(P(\cdot\mid x_{i}))-\Phi(y_{i})\right),\quad\quad A^{\prime}:=\frac{1}{k}\left(\Phi(P(\cdot\mid x_{i}))-\Phi(y_{i}^{\prime})\right),$
$B:=\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})~\text{and}~j\neq i}\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j}).$
The triangle inequality implies that
$\left\|A+B\right\|_{\mathcal{H}}\leq\|A\|_{\mathcal{H}}+\|B\|_{\mathcal{H}},\quad\quad\left\|A^{\prime}+B\right\|_{\mathcal{H}}\geq\|B\|_{\mathcal{H}}-\|A^{\prime}\|_{\mathcal{H}}.$
Therefore,
$\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y_{i}}-\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}=\|A+B\|_{\mathcal{H}}-\|A^{\prime}+B\|_{\mathcal{H}}$
$\leq\|A\|_{\mathcal{H}}+\|B\|_{\mathcal{H}}-(\|B\|_{\mathcal{H}}-\|A^{\prime}\|_{\mathcal{H}})=\|A\|_{\mathcal{H}}+\|A^{\prime}\|_{\mathcal{H}}.$
Similarly,
$\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}-\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y_{i}}\leq\|A\|_{\mathcal{H}}+\|A^{\prime}\|_{\mathcal{H}}.$
Hence,
$\left|\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y_{i}}-\left.\psi({\rm NN}(x,k,X_{n}))\right|_{y^{\prime}_{i}}\right|\leq\|A\|_{\mathcal{H}}+\|A^{\prime}\|_{\mathcal{H}}$
$=\frac{1}{k}\left\|\Phi(P(\cdot\mid x_{i}))-\Phi(y_{i})\right\|_{\mathcal{H}}+\frac{1}{k}\left\|\Phi(P(\cdot\mid x_{i}))-\Phi(y_{i}^{\prime})\right\|_{\mathcal{H}}\leq\frac{2\sqrt{2}C_{\rm ker}}{k},$
where the last inequality follows from the bound (9).
On the other hand, the output $y_{i}$ associated with any non-$k$-nearest neighbours $i\not\in{\rm NN}(x,k,X_{n})$ does not appear in $\psi({\rm NN}(x,k,X_{n}))$, so changing the value of $y_{i}$ in this case does not change $\psi({\rm NN}(x,k,X_{n}))$.
@@
The number of distinct index sets of nearest neighbours is identical to the number of distinct ways the set $X_{n}$ of $n$ points is intersected by balls $B(x,r_{k,n}(x))$ of center $x$ and radius $r_{k,n}(x)$ being the distance of the $k$-th nearest neighbour from $x$. This number is upper-bounded by the number of distinct ways $X_{n}$ is intersected by the class $\mathcal{B}=\{B(x,r)\mid x\in\mathcal{X},\ r>0\}$ of all balls, which is further upper-bounded by $n^{\mathcal{V}_{\mathcal{B}}}$ with the VC dimension $\mathcal{V}_{\mathcal{B}}$ of $\mathcal{B}$ (kpotufe2011k, p.6). Therefore, by using the union bound, the probability that the statement (23) holds for some $x$ and $k$ is upper bounded by the bound (22) times $n^{\mathcal{V}_{\mathcal{B}}}$:
@@
$\mathbb{E}\left[\left\|\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}\Phi(P(\cdot\mid x_{j}))-\Phi(y_{j})\right\|_{\mathcal{H}}^{2}\mid X_{n}\right]=\mathbb{E}\left[\left\|\frac{1}{k}\sum_{j\in{\rm NN}(x,k,X_{n})}z_{j}\right\|_{\mathcal{H}}^{2}\mid X_{n}\right]$